OpenStudy (anonymous):

Solve the system 2x + 3y = 3 and 3x – 2y = 11 by using graph paper or graphing technology. What is the solution to the system? (–3, 3) (–1, –7) (1, –4) (3, –1)

OpenStudy (ivettef365):

you can try by substituting the points on the equations to see which one is true example, lets take point (-3,3) 2x + 3y = 3 2(-3) + 3(3) = 3 -6 + 9 = 3 3 = 3 ok then 3x - 2y = 11 3(-3) - 2(3) = 11 -9 -6 = 11 -15 = 11 NOT TRUE, so this point won't satisfy both equations not good

OpenStudy (ivettef365):

or you can graph the equations, and see where they meet. example 2x + 3y = 3 3y = -2x + 3 y = -3/2x + 1

OpenStudy (anonymous):

so you would do the substitution method with the answer choices until they are both true?

OpenStudy (ivettef365):

correct

OpenStudy (anonymous):

ok then thanks